Rachael hands in her 4th exam. The scores on her first three exams were 75, 95, and 97. What does she need to earn on the 4th exam to average at least a 90 on the four exams? Construct an inequality with the variable x representing her 4th exam score, then solve the inequality for x.
Answer by rfer(16322)

(75+95+97+x)/4=90
267+x=90*4
267+x=360
x=360-267
x=93